OpenStudy (anonymous):

You have $10,000 that you plan to invest in a compound-interest bearing interest. Your investment agent advises you that you can invest the $10,000 at 8% compounded quarterly for three years or you can invest it at 8 ¼ % compounded annually for three years. Which investment should you choose to receive the most interest? Show your work for calculating the interest for both options.

OpenStudy (anonymous):

Amount earned after T years compounding C times per year with an initial investment of P and interest rat of I, is \[P(1+I/C)^{CT}\] Do you understand?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Compound_interest

OpenStudy (anonymous):

ok so the second is 10,000*1.0825=10825*1.0825=11718*1.0825=12684 the first is 10,000*1.08=10800*1.08=11664*1.08=12597*1.08 so it is already more then the other one and the first one has only been through one year so it is the 8% one